# Metric Differential Privacy at the User-Level

### v1

@@ 4.2. Unordered Release of Item-wise Queries
As PrivEMDItemWise does not hide the size of $K$, we show it satisfies bounded $d_{\textsf{EM}}$ DP. We use the following argument: for a neighboring dataset $K^{\prime}=\{x_{1}^{\prime},\ldots,x_{m}^{\prime}\}$, by Lemma 2.1 there exists a permutation $\pi:[m]\rightarrow[m]$ satisfying (3). Observe that we release the query responses in an unordered fashion by explicitly shuffling them. This allows us to pair up the element $x_{i}$ with $x_{\pi(i)}$ and analyze the privacy guarantee of releasing $\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m})$ versus $\mathcal{A}(x_{\pi(i)}^{\prime}),\mathcal{A}(x_{\pi(m)}^{\prime})$. Prior work does this with composition (Fernandes et al., 2019).
@@
In particular, our core technical contribution is to generalize the privacy amplification by shuffling to $d_{\mathcal{X}}$-DP. Specifically, we analyze the privacy guarantee between two multisets $\textsf{Shuffle}(\{\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))$ and $\textsf{Shuffle}(\mathcal{A}(x_{1}^{\prime}),\ldots,\mathcal{A}(x_{m}^{\prime}))$ when $\mathcal{A}$ satisfies $d_{\mathcal{X}}$-DP, in terms of the vector of distances $v=(d_{\mathcal{X}}(x_{i},x_{i}^{\prime}))_{i=1}^{m}$. The parameters we will be interested in are $\|v\|_{0}$, or the number of nonzero elements in $v$, and $\|v\|_{1}$ since in our different privacy models we will be able to bound both. Formally:
@@
$\Pr[\mathsf{Shuffle}(\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))=O]\\ \leq e^{\alpha}\Pr[\mathsf{Shuffle}(\mathcal{A}(x_{1}^{\prime}),\ldots,\mathcal{A}(x_{m}^{\prime}))=O]+\delta e^{\alpha},$
@@
Remarks. In particular, if $\alpha_{0}\leq\frac{\|v\|_{0}}{\|v\|_{1}}$, the above bound is roughly $\frac{\alpha_{0}\|v\|_{1}}{\sqrt{m}}$, which grows with just $\sqrt{m}$ (as $\|v\|_{1}\leq m$). The standard result is only applicable when $\mathcal{A}$ satisfies $\alpha$-local DP, and just $x_{1}$ is changed to $x_{1}^{\prime}$ (since each user owns just one item). We generalize the state-of-the-art privacy amplification by shuffling analysis of Feldman et al. (Feldman et al., 2022) to $d_{\mathcal{X}}$-DP, and our result may be of independent interest. We state our result generally in terms of the vector $\|v\|_{0},\|v\|_{1}$ since we will be applying it with different known bounds on these quantities.
Proof Sketch. We first generalize the analysis of amplification by shuffling to the datasets $(x_{1},x_{2},\ldots,x_{m})$ and $(x_{1}^{\prime},x_{2},\ldots,x_{m})$, where $d_{\mathcal{X}}(x_{1},x_{1}^{\prime})\leq w_{1}$. We show the resulting privacy parameter is given by $f(w_{1})$ where
$f(w_{1})=\ln\left(1+\frac{e^{\alpha_{0}w_{1}}-1}{e^{\alpha_{0}w_{1}}+1}\left(\frac{8\sqrt{e^{\alpha_{0}}\ln(4/\delta)}}{\sqrt{m}}+\frac{8e^{\alpha_{0}}}{m}\right)\right).$
Then, we apply group privacy $\|v\|_{0}$ times to show the general result holds with parameter $\sum_{i=1}^{\|v\|_{0}}f(\alpha_{0}w_{i})$. The function $f$ is concave so the worst-case amplification is simply $\|v\|_{0}f(\frac{\|v\|_{1}}{\|v\|_{0}})$. ∎
Comparison with Composition. Analyzing Thm. 4.3 using the state-of-the-art composition results (Kairouz et al., 2015) and $\alpha_{0}\leq 1$ gives us
@@
Combining (3) and Thm. 4.3, we obtain an improved privacy guarantee for PrivEMDItemWise. We may state this guarantee in both the bounded local and central models. In the local model, recall that each user is applying PrivEMDItemWise to their data. In the central model, the central server applies PrivEMDItemWise to the entire dataset, and thus releases the frequencies of $mn$ itemwise queries.
@@
For any $\delta\in(0,1)$, PrivEMDItemWise shown in Alg. 2 satisfies bounded local $(\alpha,\delta^{\prime})$-$d_{\textsf{EM}}$ DP, where
@@
Remarks. Thm. 4.4 gives the tightest possible privacy parameters, but we may also give an asymptotic formula as follows. For desired privacy parameters $(\alpha,\delta)$, one should set
@@
(6) $\alpha_{0}=\begin{cases}\frac{\alpha}{32\sqrt{m\ln(4me^{\alpha}/\delta)}}&\text{ if $\alpha\leq 32\sqrt{m\ln(4me^{\alpha}/\delta)}$}\\ 2\ln\left(\frac{\alpha}{16\sqrt{m\ln(4me^{\alpha}/\delta)}}\right)&32\sqrt{m\ln(4me^{\alpha}/\delta)}<\alpha<m\end{cases}$
and respectively
@@
(7) $\alpha_{0}=\begin{cases}\frac{\alpha\sqrt{n}}{32\sqrt{m\ln(4me^{\alpha}/\delta)}}&\text{ if $\alpha\sqrt{n}\leq 32\sqrt{m\ln(4me^{\alpha}/\delta)}$}\\ 2\ln\left(\frac{\alpha\sqrt{n}}{16\sqrt{m\ln(4me^{\alpha}/\delta)}}\right)&32\sqrt{m\ln(4me^{\alpha}/\delta)}<\alpha\sqrt{n}<m\sqrt{n}\end{cases}$
@@
in order to achieve $d_{\textsf{EM}}$ DP in the bounded local (respectively central) model. Assuming $\alpha\leq O(\ln(m))$, this means that the privacy parameter will be roughly $\frac{\alpha}{\sqrt{m}}$ (resp. $\ln(\frac{\alpha\sqrt{n}}{\sqrt{m}})$) for releasing the $m$ samples; this is asymptotically better than the analysis with composition which would require setting $\alpha_{0}=\frac{\alpha}{m}$ (resp. $\frac{\alpha}{m})$). Even with higher $\alpha=m^{c}$ for $c<1$, the budget is still $\frac{\alpha}{\sqrt{m^{1+c}}}$ (resp. $\ln(\frac{\alpha\sqrt{n}}{\sqrt{m^{1+c}}}$), which are both significant asymptotic improvements.
@@
Concrete Example. Our improved analysis makes the most significant improvements in the central model. Here, we would have to apply PrivEMDItemWise with $\alpha_{0}=\frac{\alpha}{m}=0.025$ for each of the $m=10^{3}$ location data points per user. Using the guarantee of Thm. 4.4, it is possible to set $\alpha_{0}\approx 3.0$ – a several orders of magnitude improvement.
5. Generalization to Unbounded DP
@@
The mechanisms presented so far face two challenges when applied to the unbounded data setting. First, a direct privacy analysis of the unbounded data setting is difficult since we cannot leverage Lemma 2.1, which significantly simplifies the analysis (for the bounded data setting). Second, and more importantly, the unbounded central model offers no utility improvement over the local model. In the worst-case scenario, a single user may contribute nearly all the data in the dataset, effectively reducing any algorithm to satisfying only local $d_{\textsf{EM}}$-DP. This issue has been noted in previous work in user-level DP (Liu et al., 2023).
@@
In this section, we tackle these challenges by showing a blackbox reduction from unbounded $d_{\textsf{EM}}$-DP to bounded $d_{\textsf{EM}}$-DP. Our reduction works in both the local and central models. The key idea of the reduction is to smoothly project a dataset $K$ of any size to a dataset $L$ of a given fixed size, such that the $d_{\textsf{EM}}$ distance between any two input datasets and the $d_{\textsf{EM}}$ distance between their projections are roughly the same. Then, it is easy to show that applying any bounded $d_{\textsf{EM}}$-DP algorithm to the smooth projections is sufficient to guarantee unbounded $d_{\textsf{EM}}$-DP for the entire scheme.
@@
Our proposed projection mechanism is smooth in a near-multiplicative sense, albeit with a small additive penalty when the $d_{\textsf{EM}}$ between the two datasets is small. We account for this subtlety, by slightly modify the privacy semantics of $d_{\textsf{EM}}$-DP in the unbounded setting to not grow arbitrarily strong as $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\rightarrow 0$. Instead, we introduce a distance threshold $r$ such that for all $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\leq r$ enjoys a uniform privacy guarantee of $\varepsilon$. This refined privacy definition, termed discrete $d_{\textsf{EM}}$-DP, is formalized (in the local model) as:
@@
$\Pr[M(K)=O]\leq e^{\varepsilon}\Pr[M(K^{\prime})=O]+\delta.$
@@
$\Pr[M(K)=O]\leq e^{\varepsilon\lceil\frac{d}{r}\rceil}\Pr[M(K^{\prime})=O]+\delta\exp(\lceil\tfrac{d}{r}\rceil).$
Fact 5.1 is implied from Definition 5.1 by a direct application of group privacy (Dwork, 2006) (proven in Lemma A.2).This guarantee can be interpreted as providing $d_{\textsf{EM}}$-DP at the granularity of units of $d_{\textsf{EM}}$ distance $r$. Note that for all $d\geq r$, we have $\varepsilon\lceil\frac{d}{r}\rceil\leq\frac{2\varepsilon}{r}d$. Thus, $(\varepsilon,\delta,r)$-discrete local $d_{\textsf{EM}}$ DP is roughly equivalent to $(\frac{2\varepsilon}{r},\delta)$-unbounded local $d_{\textsf{EM}}$-DP, except if $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\leq r$. In this case, the privacy parameter will not go below $\varepsilon$. This adjustment does not significantly alter the overall privacy semantics of $d_{\textsf{EM}}$-DP; one may simply set $\alpha$ as described in Sec. 4.
@@
$\Pr[M(K_{G})=O]\leq e^{\varepsilon}\Pr[M(K^{\prime}_{G})=O]+\delta.$
As before, $(\varepsilon,\delta,r)$-discrete central $d_{\textsf{EM}}$-DP is roughly equivalent to $(\frac{2\varepsilon}{r},\delta)$-bounded central $d_{\textsf{EM}}$-DP when all user datasets have size $m$. However, we will see that Definition 5.2 is the appropriate generalization to unbounded user datasets under our projection mechanism which is described below.
Our projection mechanism first generates a fixed number of samples with replacement from each user’s dataset $K$. Next, it applies a blackbox bounded $d_{\textsf{EM}}$-DP mechanism, $\mathcal{A}$, to the projected dataset $L$. By blackbox application we mean that $\mathcal{A}$ can be any arbitrary mechanism as long as it satisfies bounded $d_{\textsf{EM}}$-DP. The projection mechanism, in the central model, is outlined in Alg. 3 (in the local model, each user samples from their own $K_{i}$, so we would simply have $n=1$). The smoothness of the projection from $K_{G}$ to $L$ follows from the following claim:
Lemma 5.2.
Let $\tilde{K},\tilde{K}^{\prime}\in\Delta^{\mathcal{X}}$ be probability distributions, and let $C^{*}$ be the minimum cost coupling between $\tilde{K},\tilde{K}^{\prime}$. Let $\{(x_{i},y_{i})\}_{i=1}^{s}$ denote $s$ i.i.d. samples from $C^{*}$, and let $L=(x_{1},\ldots,x_{s})$ and $L^{\prime}=(y_{1},\ldots,y_{s})$. Then,
@@
Proof Sketch: Intuitively, for any coupling $C$ between $\tilde{K}_{i}$ and $\tilde{K}_{i}^{\prime}$, we can simulate sampling $s$ times from $\tilde{K}_{i}^{\prime}$ by first sampling $\{x_{1},\ldots,x_{s}\}$ from $\tilde{K}_{i}$, and then sampling $y_{i}\sim C_{x_{i}}(\cdot)$. This view shows there is a transportation plan from $L=\{x_{1},\ldots,x_{s}\}$ and $L^{\prime}=\{y_{1},\ldots,y_{s}\}$ of expected cost $\mathbb{E}_{x\sim C_{1},y\sim C_{x}}d_{\mathcal{X}}(x,y)=d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})$. Using Bernstein’s inequality, we can show with probability at least $1-\delta$, $d_{\textsf{EM}}(\tilde{L},\tilde{L}^{\prime})$ is upper bounded by $2d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})+\frac{6}{s}\log(\frac{1}{\delta})$. ∎
Thus, sampling is a smooth projection from $K_{G}$ to $L$, with the caveat that there is an additive $\frac{\log(1/\delta)}{s}$ term that comes into play if $s$ is too small to guarantee convergence. This is the reason for our relaxation to discrete $d_{\textsf{EM}}$-DP.
In summary, we have the following privacy guarantee:
@@
Let $\varepsilon>0$ and $\delta,r\in[0,1]$ be arbitrary constants. Suppose $\mathcal{M}$ is a mechanism which satisfies $(\alpha,\delta)$-bounded local $d_{\textsf{EM}}$-DP (Definition 3.1), where
@@
Then, BoundedEMDReduction satisfies $(\varepsilon,2\delta,r)$-discrete local $d_{\textsf{EM}}$-DP. Similarly, if $\mathcal{M}$ satisfies $(\alpha,\delta)$-bounded central $d_{\textsf{EM}}$-DP (Definition 3.1), then BoundedEMDReduction satisfies $(\varepsilon,2\delta,r)$-discrete central $d_{\textsf{EM}}$-DP.
Remarks. If the number of samples $s$ is at least $\frac{\ln(1/\delta)}{r}$, then Thm. 5.3 shows there is only a small multiplicative cost to considering just bounded $d_{\textsf{EM}}$-DP (in the respective local or central model). In this case, the bounded algorithm will need to roughly satisfy $(\frac{\varepsilon}{r},\delta)$-bounded $d_{\textsf{EM}}$-DP, and this is roughly the same as the resulting $(\varepsilon,\delta,r)$-discrete $d_{\textsf{EM}}$ DP algorithm. There is no privacy disadvantage to taking a large number of samples, and the utility may also increase due to more information about the dataset being captured (recall that the projection does not providing privacy; it is being provided by $\mathcal{M}$). Thus, the number of samples may be set to be large with computational costs being the only constraint.
Proof Sketch. Let $L$ and $L^{\prime}$ denote the sampled data for $K_{G}$ and $K_{G}^{\prime}$, respectively. By convexity of DP, it suffices to analyze the privacy guarantee for any coupling between the random variables $L,L^{\prime}$. In particular, we use the optimal coupling between $K_{G},K_{G}^{\prime}$ to define this coupling. The resulting privacy parameter is bounded in terms of the expected cost $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})$ by Lemma 5.2. ∎
BoundedEMDReduction can be used to bound the contribution of each user in the central setting, allowing us to apply the simpler Definition 3.2. In addition, it can be used to adapt PrivEMDItemWise to the unbounded data setting. One caveat is that utility may not be preserved if the number of user samples is too small or, in the central setting, if the users data distributions are heterogeneous. In particular, if users have varying numbers of samples, each from different distributions, applying BoundedEMDReduction equalizes the frequency of all user data. Nonetheless, it is often reasonable to assume the users have homogeneous data distributions (Liu et al., 2020; Acharya et al., 2023).
6. Applications of Proposed Mechanisms
In this section, we compare the utilities of PrivEMDLinear and PrivEMDItemWise to existing mechanisms satisfying user-level DP. For simplicity, we assume the bounded data setting.

### v2

@@ 4.2 Unordered Release of Item-wise Queries
In particular, our core technical contribution is to analyze a general formulation of the privacy amplification by shuffling problem, where the vector $x_{1},\ldots,x_{m}$ is changed to an arbitrary vector $x_{1}^{\prime},\ldots,x_{m}^{\prime}$. The key quantities we have control over are $\|v\|_{1}$ and $\|v\|_{0}$, where $v=(d_{\mathcal{X}}(x_{i},x_{i}^{\prime}))_{i=1}^{m}$ (the $d_{\textsf{EM}}$ distance allows us to bound $\|v\|_{1}$ and the maximum contribution from a user allows us to bound $\|v\|_{0}$); thus, our privacy bound depends on them. We believe our general result is of independent interest in the field of metric DP. Formally,
@@
$\Pr[\mathsf{Shuffle}(\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))=O]\leq e^{\alpha}\Pr[\mathsf{Shuffle}(\mathcal{A}(x_{1}^{\prime}),\ldots,\mathcal{A}(x_{m}^{\prime}))=O]+\delta e^{\alpha},$
@@
Remarks. In particular, if $\alpha_{0}\leq\frac{\|v\|_{0}}{\|v\|_{1}}$, the above bound is $\approx\frac{\alpha_{0}\|v\|_{1}}{\sqrt{m}}$, which grows with just $\sqrt{m}$ (as $\|v\|_{1}\leq m$). The standard shuffling result only assumes that $\mathcal{A}$ satisfies $\alpha$-local DP, and that just $x_{1}$ is changed to $x_{1}^{\prime}$ (since each user owns a single item). Theorem 4.3 can be specialized to recover the state-of-the-art result for this special case Feldman et al. (2022), but it is significantly more general in its current form.
Proof Sketch. Using group privacy, we can analyze the privacy guarantee between $(x_{1},\ldots,x_{m})$ and $(x_{1}^{\prime},\ldots,x_{m}^{\prime})$, where up to $\|v\|_{0}$ points change, instead of just changing one point as a time. We then analyze a change of one point by generalizing and simplifying the state-of-the-art technique in Feldman et al. (2022). We show that the resulting privacy parameter for changing the point $x_{i}$ to $x_{i}^{\prime}$ is $g(w_{i})$ where
$g(w_{i})=\ln\left(1+\frac{e^{\alpha_{0}w_{i}}-1}{e^{\alpha_{0}w_{i}}+1}\left(\frac{8\sqrt{e^{\alpha_{0}}\ln(4/\delta)}}{\sqrt{m}}+\frac{8e^{\alpha_{0}}}{m}\right)\right).$
By group privacy $\|v\|_{0}$ times, the overall privacy parameter is $\sum_{i=1}^{\|v\|_{0}}g(w_{i})$. The final step is proving that $g$ is concave so the worst-case amplification is simply $\|v\|_{0}g(\frac{\|v\|_{1}}{\|v\|_{0}})$. ∎
Comparison with Composition. Analyzing Theorem 4.3 using the state-of-the-art composition results (Kairouz et al., 2015) and $\alpha_{0}\leq 1$ gives us
@@
Combining (3) and Theorem 4.3, we obtain an improved privacy guarantee for PrivEMDItemWise. The guarantee can be stated in both the bounded local and central models. In the local model, recall that each user is applying PrivEMDItemWise to their data. In the central model, the aggregator applies PrivEMDItemWise to the entire dataset, and releases the frequencies of $mn$ itemwise queries.
@@
For any $\delta\in(0,1)$, PrivEMDItemWise shown in Algorithm 2 satisfies bounded local $(\alpha,\delta^{\prime})$-$d_{\textsf{EM}}$ DP, where
@@
Remarks. Theorem 4.4 gives the tightest possible privacy parameters, but we may also give an asymptotic formula as follows. For desired privacy parameters $(\alpha,\delta)$, one should set
@@
$\alpha_{0}=\begin{cases}\frac{\alpha}{32\sqrt{m\ln(4me^{\alpha}/\delta)}}&\text{ if $\alpha\leq 32\sqrt{m\ln(4me^{\alpha}/\delta)}$}\\ 2\ln\left(\frac{\alpha}{16\sqrt{m\ln(4me^{\alpha}/\delta)}}\right)&32\sqrt{m\ln(4me^{\alpha}/\delta)}<\alpha<m\end{cases}$ (6)
and
@@
$\alpha_{0}=\begin{cases}\frac{\alpha\sqrt{n}}{32\sqrt{m\ln(4me^{\alpha}/\delta)}}&\text{ if $\alpha\sqrt{n}\leq 32\sqrt{m\ln(4me^{\alpha}/\delta)}$}\\ 2\ln\left(\frac{\alpha\sqrt{n}}{16\sqrt{m\ln(4me^{\alpha}/\delta)}}\right)&32\sqrt{m\ln(4me^{\alpha}/\delta)}<\alpha\sqrt{n}<m\sqrt{n}\end{cases}$ (7)
@@
in order to achieve $d_{\textsf{EM}}$-DP in the bounded local and central model, respectively. Assuming $\alpha\leq O(\ln(m))$, this means that the privacy parameter will be roughly $\frac{\alpha}{\sqrt{m}}$ (resp. $\ln(\frac{\alpha\sqrt{n}}{\sqrt{m}})$) for releasing the $m$ samples; this is asymptotically better than the analysis with composition which would require setting $\alpha_{0}=\frac{\alpha}{m}$ (resp. $\frac{\alpha}{m})$). Even with higher $\alpha=m^{c}$ for $c<1$, the budget is still $\frac{\alpha}{\sqrt{m^{1+c}}}$ (resp. $\ln(\frac{\alpha\sqrt{n}}{\sqrt{m^{1+c}}}$)), which are both significant asymptotic improvements.
@@
Concrete Example. Our improved analysis makes the most significant improvements in the central model. Here, we would have to apply PrivEMDItemWise with $\alpha_{0}=\frac{\alpha}{m}=0.025$ for each of the $m=10^{3}$ location data points per user. Using the guarantee of Theorem 4.4, it is possible to set $\alpha_{0}\approx 3.0$ – a several orders of magnitude improvement.
Open Questions. The primary open question in the design of our mechanism is whether one can obtain a tighter privacy analysis of Theorem 4.4 that does not rely on group privacy, but rather analyzes the amplification potential of all points at once.
New Proof Techniques. First, in the case of linear queries, we bound the sensitivity under $d_{\textsf{EM}}$-DP using a Lipschitz assumption. This stands in contrast to sensitivity arguments in standard DP, which typically assume the weaker condition that the query is merely bounded. This is a novel aspect of our analysis that has not been explored in prior sensitivity analysis. Second, in the context of unordered release of item-wise queries, we prove a new result for privacy amplification by shuffling. This is a more generalized version of the standard shuffling result Feldman et al. (2022) in two ways – $(1)$ the standard result assumes that the private mechanism satisfies local-DP whereas we work with mechanisms that satisfy $d_{\mathcal{X}}$-DP; $(2)$ the standard result only considers changes to a single point, whereas we allow changes to all $m$ input points. Using this result, we provide a tighter privacy analysis than what composition theorems would allow. Our new privacy amplification by shuffling result can be of independent interest in the field of metric DP.
5 Generalization to Unbounded DP
@@
The mechanisms presented so far face two challenges when applied to the unbounded data setting. First, a direct privacy analysis of the unbounded data setting is difficult since we cannot leverage Lemma 2.1, which significantly simplifies the analysis (for the bounded data setting). Second, and more importantly, the unbounded central model offers no utility improvement over the local model. In the worst-case scenario, a single user may contribute nearly all the data in the dataset, effectively reducing any algorithm to satisfying only local $d_{\textsf{EM}}$-DP. This issue has been noted in previous work in user-level DP Liu et al. (2023).
@@
In this section, we tackle these challenges by showing a blackbox reduction from unbounded $d_{\textsf{EM}}$-DP to bounded $d_{\textsf{EM}}$-DP. Our reduction works in both the local and central models. The key idea of the reduction is to smoothly project a dataset $K$ of any size to a dataset $L$ of a given fixed size, such that the $d_{\textsf{EM}}$ distance between any two input datasets and the $d_{\textsf{EM}}$ distance between their projections are roughly the same. Then, it is easy to show that applying any bounded $d_{\textsf{EM}}$-DP algorithm to the smooth projections is sufficient to guarantee unbounded $d_{\textsf{EM}}$-DP for the entire scheme. Full proofs for this section appear in Appendix C.
@@
Our proposed projection mechanism is smooth in a near-multiplicative sense, albeit with a small additive penalty when the $d_{\textsf{EM}}$ between the two datasets is small. We account for this subtlety by slightly modifying the privacy semantics of $d_{\textsf{EM}}$-DP in the unbounded setting to not grow arbitrarily strong as $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\rightarrow 0$. Instead, we introduce a distance threshold $r$ such that all $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\leq r$ enjoys a uniform privacy guarantee of $\varepsilon$. This refined privacy definition, termed discrete $d_{\textsf{EM}}$-DP, is formalized (in the local model) as:
@@
$\Pr[\mathcal{M}(K)=O]\leq e^{\varepsilon}\Pr[\mathcal{M}(K^{\prime})=O]+\delta.$
@@
$\Pr[\mathcal{M}(K)=O]\leq e^{\varepsilon\lceil\frac{d}{r}\rceil}\Pr[\mathcal{M}(K^{\prime})=O]+\delta\exp(\lceil\tfrac{d}{r}\rceil).$
Fact 5.1 is implied from Definition 5.1 followed by a direct application of group privacy Dwork (2006). This guarantee can be interpreted as providing $d_{\textsf{EM}}$-DP at the granularity of units of $d_{\textsf{EM}}$ distance $r$. Note that for all $d\geq r$, we have $\varepsilon\lceil\frac{d}{r}\rceil\leq\frac{2\varepsilon}{r}d$. Thus, $(\varepsilon,\delta,r)$-discrete local $d_{\textsf{EM}}$ DP is roughly equivalent to $(\frac{2\varepsilon}{r},\delta)$-unbounded local $d_{\textsf{EM}}$-DP, except if $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})\leq r$. In this case, the privacy parameter will not go below $\varepsilon$. This adjustment does not significantly alter the overall privacy semantics of $d_{\textsf{EM}}$-DP; one may simply set $\alpha$ as described in Section 4.
@@
$\Pr[\mathcal{M}(K_{G})=O]\leq e^{\varepsilon}\Pr[\mathcal{M}(K^{\prime}_{G})=O]+\delta.$
As before, $(\varepsilon,\delta,r)$-discrete central $d_{\textsf{EM}}$-DP is roughly equivalent to $(\frac{2\varepsilon}{r},\delta)$-bounded central $d_{\textsf{EM}}$-DP when all user datasets have size $m$. We will see that Definition 5.2 is the appropriate generalization to unbounded user datasets under our projection mechanism which is described below.
Because our projection mechanism must preserve the $d_{\textsf{EM}}$ between $\tilde{K},\tilde{K}^{\prime}$, it is not acceptable to select an arbitrary set of points from $\tilde{K},\tilde{K}^{\prime}$, as this could dramatically inflate the $d_{\textsf{EM}}$ distance. However, couplings are intimately tied with sampling – observe that two random samples from $\tilde{K},\tilde{K}^{\prime}$ can actually be coupled so that their expected distance is $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})$. If this process is repeated multiple times, then the cost of coupling the samples will converge to $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})$. Thus, sampling with replacement will result in a projection that does not increase the $d_{\textsf{EM}}$ distance by too much.
@@
Lemma 5.2.
Let $\tilde{K},\tilde{K}^{\prime}\in\Delta^{\mathcal{X}}$ be probability distributions, and let $C^{*}$ be the minimum cost coupling between $\tilde{K},\tilde{K}^{\prime}$. Let $\{(x_{i},y_{i})\}_{i=1}^{s}$ be $s$ i.i.d. samples from $C^{*}$, $L=(x_{1},\ldots,x_{s})$ and $L^{\prime}=(y_{1},\ldots,y_{s})$. Then,
Remarks. The multiplicative factor of $1+\sqrt{2}$ shows that the projection is smooth when $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})$ dominates the additive factor of $\frac{3}{s}\ln(\frac{1}{\delta})$. This is achieved when the number of samples $s$ is much larger than the inverse of $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})$. If $s$ is not sufficiently large, then not enough samples are being taken to ensure convergence.
Proof Sketch. Let $C$ be a minimum-cost coupling between $\tilde{K},\tilde{K}^{\prime}$. To show that $\tilde{L},\tilde{L}^{\prime}$ are not far from each other, we can view $\tilde{L}^{\prime}$ as being generated from $\tilde{L}$, where for each point $x\in L$, a point $y\sim C_{x}(\cdot)$ is added to $L^{\prime}$. This view of $\tilde{L},\tilde{L}^{\prime}$ shows there is a transportation plan from $L=\{x_{1},\ldots,x_{s}\}$ and $L^{\prime}=\{y_{1},\ldots,y_{s}\}$ of expected cost $\mathbb{E}_{x\sim C_{1},y\sim C_{x}}d_{\mathcal{X}}(x,y)=d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})$. Using Bernstein’s inequality, we can show with probability at least $1-\delta$, $d_{\textsf{EM}}(\tilde{L},\tilde{L}^{\prime})$ is upper bounded by $2d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})+\frac{6}{s}\log(\frac{1}{\delta})$. ∎
Our full reduction to bounded $d_{\textsf{EM}}$-DP first projects $K$ onto a dataset of size $m$ by taking samples with replacement. Next, it applies a blackbox bounded $d_{\textsf{EM}}$-DP mechanism, $\mathcal{A}$, to the projected dataset $L$. By blackbox application we mean that $\mathcal{A}$ can be any arbitrary mechanism as long as it satisfies bounded $d_{\textsf{EM}}$-DP. We call this mechanism BoundedEMDReduction, and it is illustrated in the central model in Algorithm 3 (in the local model, each user samples from their own $K_{i}$, so we would simply have $n=1$). Using the projection guarantee of Lemma 5.2, BoundedEMDReduction enjoys the following privacy guarantee:
@@
Let $\varepsilon>0$ and $\delta,r\in[0,1]$ be arbitrary constants. Suppose $\mathcal{A}$ is a mechanism which satisfies $(\alpha,\delta)$-bounded local $d_{\textsf{EM}}$-DP (Definition 3.1), where
@@
Then, BoundedEMDReduction satisfies $(\varepsilon,2\delta,r)$-discrete local $d_{\textsf{EM}}$-DP. Similarly, if $\mathcal{A}$ is $(\alpha,\delta)$-bounded central $d_{\textsf{EM}}$-DP (Definition 3.1), then BoundedEMDReduction is $(\varepsilon,2\delta,r)$-discrete central $d_{\textsf{EM}}$-DP.
Remarks. If the number of samples $s$ is at least $\frac{\ln(1/\delta)}{r}$, then Theorem 5.3 shows there is only a small multiplicative cost to considering just bounded $d_{\textsf{EM}}$-DP (in the respective local or central model). In this case, $\mathcal{A}$ will need to roughly satisfy $(\frac{\varepsilon}{r},\delta)$-bounded $d_{\textsf{EM}}$-DP, and this is roughly the same as the resulting $(\varepsilon,\delta,r)$-discrete $d_{\textsf{EM}}$ DP algorithm. There is no privacy disadvantage to taking a large number of samples, and the utility may also increase due to more information about the dataset being captured (recall that the projection does not providing privacy; it is being provided by $\mathcal{A}$). Thus, the number of samples may be set to be large with computational costs being the only constraint.
Proof Sketch. The proof of privacy is almost immediate from Lemma 5.2. The variables $\tilde{L},\tilde{L}^{\prime}$ in two executions of BoundedEMDReduction are random variables, but with probability $\delta$, the $d_{\textsf{EM}}$ distance between them is $c=(1+\sqrt{2})r+\frac{3}{s}\ln(\frac{1}{\delta})$. By the convexity of differential privacy, we can analyze the privacy parameter of every fixed choice of $\tilde{L},\tilde{L}^{\prime}$. With probability $1-\delta$, the privacy parameter will be $\frac{\alpha}{c}$, by the privacy guarantee of $\mathcal{A}$. ∎
BoundedEMDReduction can be used to bound the contribution of each user in the central setting, allowing us to apply the simpler Definition 3.2. In addition, it can be used to adapt PrivEMDItemWise to the unbounded data setting. One caveat is that utility may not be preserved if the number of user samples is too small or, in the central setting, if the users data distributions are heterogeneous. In particular, if users have varying numbers of samples, each from different distributions, applying BoundedEMDReduction equalizes the frequency of all user data. Nonetheless, it is often reasonable to assume the users have homogeneous data distributions Liu et al. (2020); Acharya et al. (2023).
Open Questions. Many sampling procedures are likely to be compatible with $d_{\textsf{EM}}$. This leads to the question of whether different procedures, such as sampling without replacement, are also smooth projections.
New Proof Techniques. Our blackbox reduction from unbounded to bounded $d_{\textsf{EM}}$-DP involves projecting a dataset onto one with a fixed size. The projection needs to be smooth, i.e., the $d_{\textsf{EM}}$ distance between any two input datasets and the $d_{\textsf{EM}}$ distance between their projections are roughly the same. We prove that sampling with replacement is indeed a smooth projection. The novelty of the analysis comes from viewing sampling from two datasets in terms of a coupling between them.
6 Applications of Proposed Mechanisms
In this section, we compare the utilities of PrivEMDLinear and PrivEMDItemWise to existing mechanisms satisfying user-level DP. For simplicity, we assume the bounded data setting. Full proofs for this section appear in Appendix D.